OpenStudy (anonymous):

-5^3sqrt625 + ^3sqrt40

OpenStudy (radar):

\[-5^{3}\sqrt{625}+3\sqrt{40}\]

jhonyy9 (jhonyy9):

,,radar" i think this ^3 wann to be cuberoot sign ...???

OpenStudy (radar):

Is the -5 with an exponent of cube root of 625?

jhonyy9 (jhonyy9):

no there is cuberoot of 625 and cuberoot of 40 -- i think

OpenStudy (radar):

\[-5^{\sqrt[3]{625}}?????\]

jhonyy9 (jhonyy9):

,,jmallis" -so how do you think ,please ?

jhonyy9 (jhonyy9):

,,jmallis" -so how do you think ,please ?

OpenStudy (radar):

\[-5 \sqrt[3]{625}+\sqrt[3]{40}\]

OpenStudy (anonymous):

right the second time with the 3

jhonyy9 (jhonyy9):

,,radar" exactly like ...

jhonyy9 (jhonyy9):

625=(5^3)*5 40=(2^3)*5

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (radar):

\[-5(5)\sqrt[3]{5}+2\sqrt[3]{5}\]

OpenStudy (radar):

\[-25\sqrt[3]{5}+2\sqrt[3]{5}=-23\sqrt[3]{5}\]

OpenStudy (radar):

Not sure if this is the solution as I am not sure I have translated the problem. I have my doubts about the -5

OpenStudy (radar):

@jhonyy9 do you think that is it? The -5 casts a doubt.

OpenStudy (radar):

Anyway with what you have provided should be enough to finish it.

OpenStudy (anonymous):

ok thanks